# TeleRAG: Efficient Retrieval-Augmented Generation Inference with Lookahead Retrieval

### v2

@@ 5.2. Experiment Setups
Hardware setups. We evaluate TeleRAG on two hardware environments, Desktop and Server, which are equipped to represent the settings for the desktop and data center use cases. The Desktop has the RTX4090 (24 GB memory), and we use 3B and 8B models. The Server has the H100 (80 GB memory), and we use 8B and 13B models. These sizes represent common model sizes for RAG applications (Asai et al., 2023, 2024a). We do not evaluate the 13B model on Desktop as its model size (26 GB) exceeds RTX4090’s memory capacity. Table 3 summarizes the hardware configurations.
@@
Nprobe and top-$k$. For the IVF index, nprobe is a hyperparameter that controls the trade-off between search efficiency and retrieval quality. A common heuristic is to set nprobe to $4\sqrt{N_{c}}$ (Zilliz, 2020). Given our index size of $N_{c}=4096$, we use nprobe = 256 ($4\sqrt{4096}$) by default, unless otherwise specified. For retrieval, top-$k$ denotes the number of most relevant documents to return. We use top-$k$ = 3, which is a standard choice in RAG.
@@
RAG pipeline implementation. We implement the RAG pipelines with the FlashRAG framework (Jin et al., 2024b). For IRG, FLARE, and S-RAG, we use the framework’s default implementations. For the other pipelines, we reimplement them using FlashRAG’s APIs.
@@
Benchmark methodology. We follow a benchmarking methodology from SGLang (Zheng et al., 2024) and use GPT-3.5-Turbo (OpenAI, [n. d.]) to run through each pipeline and record the input and output text of each step. During latency evaluation, we set the LLM’s output length based on the recorded text. This way, we ensure a fair latency comparison across different LLM models.
@@
Figure 8. End-to-end latency speedup of TeleRAG and baseline on six RAG pipelines and three datasets, with RTX4090 GPU. nprobe is 256.
@@
Baseline systems. To evaluate the latency of each pipeline, we construct a clean execution flow in Python that only contains LLM generation, datastore retrieval, and other necessary logical operations to fulfill each pipeline. For LLM generation, we use SGLang (Zheng et al., 2024), which is a state-of-the-art LLM inference engine. For retrieval, we use a popular retrieval library, Faiss (Douze et al., 2024), as the CPU-offloaded baseline. [^6] This baseline resembles frameworks like LlamaIndex and LangChain, which primarily function as wrappers around LLM serving and retrieval libraries. We build a system from the ground up to avoid the overhead introduced by their complex layers designed for advanced functionality.
@@
Prefetching budget setups. Based on the methodology we described in §4.2, we profile each RAG pipeline with a small amount of 64 random samples from NQ (Kwiatkowski et al., 2019) and derive the prefetching budget of each pipeline.
(a) End-to-end throughput with six pipelines and different batch sizes using Llama-3-8B.
(b) End-to-end throughput with six pipelines and different batch sizes using Llama-2-13B.
@@
Figure 9. End-to-end throughput of TeleRAG and the baseline across six RAG pipelines on the NQ dataset using Llama-3-8B and Llama-2-13B at different batch sizes on a H100 GPU. The nprobe is set to 256, and the $x$-axis represents the batch size.
@@
Max prefetching memory. We set a maximum GPU memory limit for prefetching in each configuration. For Server, we allocate up to 16 GB. For Desktop, we allocate up to 10 GB and 3.75 GB for the 3B and 8B models, respectively. These settings demonstrate that TeleRAG can efficiently operate using only a small fraction (up to 40% for RTX4090 and 20% for H100) of total GPU memory.
5.3. Evaluation Results
@@
Single-query latency on RTX4090. We evaluate the end-to-end RAG latency for single queries on Desktop equipped with RTX4090, representing a typical local scenario. Figure 8 shows the latency reduction of TeleRAG across three datasets and two LLMs (Llama-3.2-3B and Llama-3-8B).
As shown in Figure 8, TeleRAG consistently outperforms the baseline across all tested configurations. With Llama-3.2-3B, TeleRAG achieves average speedups of $1.55\times$, $1.54\times$, and $1.49\times$ on NQ, HotpotQA, and TriviaQA, respectively. The best speedup of $2.11\times$ is achieved in the Iter-RetGen pipeline on HotpotQA, because this pipeline involves frequent retrieval operations and generally short LLM outputs, which enhances the relative impact of retrieval acceleration.
@@
Another notable improvement is in the SubQuestion pipeline, where TeleRAG achieves approximately $1.85\times$ speedup across all datasets. This pipeline uses LLM-generated sub-questions and performs batched retrievals of 3 to 4 queries simultaneously. CPU-based retrieval suffers from limited parallelism in such scenarios, but TeleRAG efficiently utilizes GPU parallelism, significantly enhancing performance.
@@
Figure 10. Simulated throughput [^8] Results for 1 GPU are real, not simulated.of TeleRAG on NQ dataset with different number of H100 GPUs.
Multi-query throughput on H100. To evaluate the TeleRAG’s performance in batched inference, we evaluate the end-to-end throughput on Server (H100) in batch sizes 1, 2, 4, and 8. We show the results of six evaluated RAG pipelines with Llama-3-8B and Llama-2-13B in Figure 9.
[…]
Figure 11. Latency breakdown for Llama-3-8B on NQ with a H100 GPU in different batch sizes. nprobe is 256.
Figure 12. Comparison of end‑to‑end retrieval latency for two batching strategies: naive mini‑batching and similarity‑aware greedy grouping. The Greedy Group bars are too short to be visible.
Multi-GPU Throughput. We also evaluate TeleRAG in a multi-GPU system to further show the scalability. For the evaluation, we construct a simulation framework around our single-GPU implementation. We first define a global batch size that the whole system will process at once, and then we construct mini-batches to distribute among each GPU. To form a mini-batch, we perform a greedy search based on the embedding similarity of queries, grouping similar queries into the same mini-batch to maximize the cluster hit rate. We present the simulated throughput of TeleRAG with a global batch size of 128 and mini-batch size of 4 with 1 to 8 GPUs. As shown in Figure 8, TeleRAG’s throughput scales well with the number of GPUs. When compared to 1 GPU results, the averaged throughput enhancements of TeleRAG are $1.7\times$, $3.1\times$, and $5.4\times$ with $2$, 44, and $8$ GPUs, respectively.
5.4. Analysis and Sensitivity Study
Latency breakdown. We further show the latency breakdown of running RAG pipelines with Llama-3-8B and Llama-2-13B on a single H100 GPU in different batch sizes in Figure 11. From Figure 11, we can observe that LLM latency grows sub-linearly with larger batch sizes. However, the latency for Faiss retrieval on CPU grows linearly with the batch size, dominating the overall latency when batch size is large. These results echo our findings in Figure 9, and show limited scalability of CPU retrieval in serving scenario. In contrast, TeleRAG significantly accelerates across all batch sizes and achieves a higher speedup from $1.4\times$ to $1.8\times$ when the batch size increases from 1 to 8.
Ablation on mini-batch strategies. In our multi-GPU results, we implement a greedy grouping strategy for mini-batching based on queries’ similarity in a global batch. Here, we examine the benefits and overhead of this strategy in Figure 12, where we show the end-to-end retrieval latency for mini-batch size $4$ with four GPUs and a global batch of 128 queries. The grouping overhead is minimal and the overall retrieval latency of greedy grouping mini-batching consistently outperform naive mini-batching across all pipelines.
@@ 5.4. Analysis and Sensitivity Study
Figure 13. Single-query retrieval speedup on NQ with different nprobe values.
@@
Table 4. The prefetch budget and corresponding averaged cluster hit rate for each pipeline and hardware setup on NQ dataset. The target retrieval nprobe is 256.
@@
Prefetch budgets and cluster hit rates. Table 4 shows the prefetch budgets we set with the profile-guided approach on RTX4090 and H100 for NQ. It also presents the average cluster hit rate achieved with this prefetch budget. From the table, we can see that it generally achieves a high cluster hit rate (¿5050%) when TeleRAG has a large prefetching budget. For cases where the budget is less than 2 GB, we observe a relatively lower hit rate (¡5050%), limiting the benefits of reducing the CPU’s search workloads. However, as observed from Figure 8, TeleRAG achieves from $1.2\times$ to $1.6\times$ end-to-end speedups for these pipelines, thanks to the combined benefit of reducing CPU workload and utilizing the GPU to perform sorting on similarity distances.
6. Related Work
6.1. Systems for RAG
RAGCache (Jin et al., 2024a) proposes a caching system that stores KV caches from datastore chunks in an order-aware manner to improve the time to the first token. This approach only reduces prefill latency, leaving retrieval and decode times unaffected, despite the fact they typically dominate total latency (Agrawal et al., 2024). Moreover, it assumes repeated use of the same document across multiple requests, limiting its scalability over large data stores. Similarly, TurboRAG (Lu et al., 2024) precomputes KV caches from the data store, but it also only optimizes prefill latency. CacheBlend introduces a selective KV-cache fusion technique for RAG to reuse pre-computed caches (Yao et al., 2024).
RaLMSpec (Zhang et al., 2024b) proposes speculative retrieval and batched verification, Chameleon (Jiang et al., 2023d) proposes a CPU-GPU-FPGA heterogeneous architecture for accelerating the retrieval process, and PipeRAG (Jiang et al., 2024) is an algorithm-system co-design technique to overlap the retrieval and generation by modifying the algorithms. However, these works focus on the paradigm of RAG that retrieves documents once every few tokens and do not apply to modular RAG applications that are widely used now, as discussed in §2.1.
Speculative RAG (Wang et al., 2024b) introduces drafting by smaller LLMs to reduce RAG latency, but it does not target the retrieval latency. APIServe (InferCept) (Abhyankar et al., 2024) proposes a novel KV cache management strategy that can support the interception of LLM generation by other workloads, including retrieval. However, this work again does not focus on optimizing retrieval latency. EdgeRAG (Seemakhupt et al., 2024) reduces the memory requirement of retrieval by re-generating and caching document embeddings at runtime, targeting extremely resource-constrained environments. RAGO (Jiang et al., 2025) introduces an abstraction for RAG to automatically pick task placement, resource allocation, and batching policies.
Unlike all these prior works, we tackle system challenges of long retrieval latency and large memory requirement for modular RAG pipelines with large-scale datastores.
6.2. Systems for Compound LLM Applications
@@
Apart from RAG, there is a growing interest in compound or agentic LLM applications, where multiple LLM calls and other applications are combined to serve complex functionalities (Zaharia et al., 2024; Berger and Zorn, 2024; Wang et al., 2024a). LLMCompiler (Kim et al., 2023a) is a framework that optimizes the execution of multiple functions in large language models by enabling parallel function calling. AI Metropolis (Xie et al., 2024) accelerates LLM-based multi-agent simulations with out-of-order execution. RAG is a specific type of application in this broader direction, and we propose systems techniques to optimize its execution latency, focusing on the characteristics of retrieval workload.
6.3. Vector Index
Vector index is a key component of RAG (Briggs et al., [n. d.]; Gao et al., 2023b), and many works have been proposed to improve their efficiency. ScaNN proposes an anisotropic quantization method for better efficiency (Guo et al., 2020). DiskANN and SPANN propose memory-disk hybrid indexing systems that work beyond the limitation of memory capacity (Chen et al., 2021; Jayaram Subramanya et al., 2019). Other prior works propose hardware acceleration of vector index with GPUs (Johnson et al., 2019), FPGAs (Zeng et al., 2022; Jiang et al., 2023a), TPU (Chern et al., 2022), or ray tracing hardware (Zhu, 2022; Liu et al., 2023). BANG proposes a method to scale graph-based ANN beyond GPU memory (Khan et al., 2024), and Rummy allows the index to scale beyond GPU memory capacity with reordered pipelining (Zhang et al., 2024a). These methods either require algorithm modifications or are bottlenecked by CPU-GPU bandwidth. Our proposal, focuses on the context of modular RAG applications, where queries for retrieval are usually generated by LLMs, and optimizes the latency and the GPU memory consumption without altering the algorithm of the IVF index.
7. Conclusion
In this paper, we introduced TeleRAG, an inference system that tackles the system challenges of RAG pipelins under latency-sensitive scenarios. By using lookahead retrieval, which overlaps data transfer and GPU computation for faster retrieval, a profile-guided approach to determine optimal prefetching data amount, and GPU-CPU cooperation, TeleRAG speeds up the end-to-end latency with minimal GPU memory requirement. Our evaluation shows that TeleRAG significantly improves performance compared to existing state-of-the-art solutions.

### v3

@@ 5.2 Experiment Setups
Hardware setups. We evaluated TeleRAG on three hardware environments, Desktop, Server1 and Server2, which are equipped to represent the settings for the desktop and data center use cases. The Desktop has the RTX4090, and we use the 3B and 8B models, which are the commonly used model sizes for RAG applications asai2023self; openscholar. For Server1 (H100) and Server2 (H200), we use the 8B model. We use Server1 for single GPU evaluation and Server2 (8 GPUs) for multi-GPU evaluation. Table 2 summarizes the hardware configurations.
@@
Nprobe and top-$k$. A common heuristic for the IVF index is to set nprobe to $4\sqrt{N_{c}}$ zilliz-blog. Given our index size of $N_{c}=4096$, we use nprobe = 256 ($4\sqrt{4096}$) by default, unless otherwise specified. For retrieval, we use top-$k$ = 3 for the number of documents to return, unless otherwise specified.
@@
RAG pipeline implementation. We implemented the RAG pipelines with the FlashRAG framework jin2024flashrag. For IRG, FLARE, and S-RAG, we used the framework’s default implementations. For the other pipelines, we reimplemented them using FlashRAG’s APIs.
@@
Benchmark methodology. We follow a benchmarking methodology from SGLang zheng2024_sglang and use GPT-3.5-Turbo gpt3-5-turbo to run through each pipeline and record the input and output text of each step. During latency evaluation, we set the LLM’s output length based on the recorded text. This way, we ensure a fair latency comparison across different LLM models.
@@
Figure 9: End-to-end latency speedup of TeleRAG and baseline on six RAG pipelines and three datasets, with RTX4090 GPU.
@@
Baseline systems. To evaluate the latency of each pipeline, we constructed a clean execution flow in Python that only contains LLM generation, datastore retrieval, and other necessary logical operations to fulfill each pipeline. For LLM generation, we used SGLang zheng2024_sglang, which is a state-of-the-art LLM inference engine. For retrieval, we used a popular retrieval library, Faiss douze2024faiss, as the CPU-offloaded baseline.
@@
Prefetching budget setups. Based on the methodology we described in §4.1, we profiled each RAG pipeline with 64 random samples from NQ kwiatkowski2019natural and derived the prefetching budget of each pipeline. Since H100 and H200 have the same CPU–GPU bandwidth, we set the same prefetch budget for Server1 and Server2.
@@
Figure 10: End-to-end throughput of TeleRAG and the CPU-offload baseline across six RAG pipelines on the NQ dataset using Llama-3-8B, evaluated at different batch sizes on a H100 GPU.
(a) End-to-end throughput on NQ.
(b) End-to-end throughput on HotpotQA.
(c) End-to-end throughput on TriviaQA.
Figure 11: Throughput scaling of TeleRAG across multiple H200 GPUs. Global batch fixed at 128; micro-batch fixed at 4.
@@
Max GPU memory for retrieval. We set a maximum GPU memory limit for prefetching in each configuration. For Server1 and Server2, we allocated 12 GB and 24 GB, respectively. For Desktop, we allocated up to 10 GB and 3.75 GB for the 3B and 8B models, respectively.
Cache Setup. The memory allocated for retrieval is shared between prefetching and caching. In our evaluation, we set the cache proportion to 50% (i.e., cache can use at most half of this memory). Since the benefit of caching is negligible on a single GPU (demonstrated in the ablation study), we enabled it only for the multi-GPU experiments on Server2.
5.3 Evaluation Results
@@
Single-query latency on RTX4090. We evaluated the end-to-end RAG latency for single query on Desktop equipped with a RTX4090, representing the typical local scenario. Figure 9 shows the latency reduction of TeleRAG across three datasets and two LLMs (Llama-3.2-3B and Llama-3-8B). As Figure 9 shows, TeleRAG consistently outperforms the CPU-offload baseline across all evaluated configurations. With Llama-3.2-3B, TeleRAG achieves average speedups of $1.55\times$, $1.54\times$, and $1.49\times$ on NQ, HotpotQA, and TriviaQA, respectively.
@@
Among the test pipelines, the best speedup of $2.11\times$ is achieved in the Iter-RetGen pipeline on HotpotQA, because this pipeline involves frequent retrieval operations and generally short LLM outputs, which enhances the relative impact of retrieval acceleration. Another notable improvement is in the SubQuestion pipeline, where TeleRAG achieves approximately $1.85\times$ speedup across all datasets. This pipeline uses LLM-generated sub-questions and performs batched retrievals of 3 to 4 queries in its retrieval stage. CPU-based retrieval suffers from limited parallelism in such scenarios, but TeleRAG efficiently utilizes GPU parallelism, significantly enhancing performance.
@@
Figure 12: Latency breakdown for Llama-3-8B on NQ with a H100 GPU in different batch sizes. nprobe is 256.
Figure 13: Throughput of TeleRAG on the NQ dataset with different numbers of H200 GPUs (with and without cache).
@@
Figure 14: Single-query retrieval speedup on NQ with different nprobe values.
Figure 15: Comparison of end-to-end latency for prefetching and cache-aware schedulers on 4 H200 GPUs. The overhead bars of prefetch schedule only are too short to be visible.
@@
Table 3: The prefetch budget and corresponding averaged cluster hit rate for each pipeline and hardware setup on NQ dataset. The target retrieval nprobe is 256.
Multi-query throughput on H100. To illustrate TeleRAG’s performance in batched inference, we evaluated the end-to-end throughput on Server1 (H100) using batch sizes 1, 2, 4, and 8. The results of six RAG pipelines with Llama-3-8B are presented in Figure 10.
[…]
Multi-GPU throughput. We evaluate TeleRAG’s scalability on Server2 (H200), a multi-GPU system. Figure 11 reports throughput for a global batch size of 128 with a micro-batch size of 4 across 1–8 GPUs, with both the prefetching and cache schedulers enabled. TeleRAG scales well with the number of GPUs: on NQ, compared to single GPU case, the average speedups are $2.0\times$, $3.8\times$, and $6.5\times$ on 2, 4, and 8 GPUs, respectively. The mild sub-linear scaling at higher GPU counts is attributed to execution time variance across micro-batches—specifically, a long tail of higher-latency batches that create load imbalance. These results still demonstrate that the lookahead retrieval technique in TeleRAG can scale effectively.
5.4 Analysis and Sensitivity Study
Latency breakdown. We further show the latency breakdown of running RAG pipelines with Llama-3-8B on a single H100 GPU in different batch sizes in Figure 12. From Figure 12, we can observe that LLM latency grows sub-linearly with larger batch sizes. However, the latency for Faiss retrieval on CPU grows linearly with the batch size, dominating the overall latency when batch size is large. These results echo our findings in Figure 10, and show limited scalability of CPU retrieval in serving scenarios. In contrast, TeleRAG significantly accelerates across all batch sizes and achieves a higher speedup from $1.4\times$ to $1.8\times$ when the batch size increases from 1 to 8.
@@ 5.4 Analysis and Sensitivity Study
Prefetch budgets and cluster hit rates. Table 3 shows the prefetch budgets we set with the profile-guided approach on RTX4090 and H100 for NQ. It also presents the average cluster hit rate achieved with this prefetch budget. From the table, we can see that it generally achieves a high cluster hit rate ($50\%$}) when TeleRAG has a large prefetching budget. For cases where the budget is less than 2 GB, we observe a relatively lower hit rate ($50\%$}), limiting the benefits of reducing the CPU’s search workloads. However, as observed from Figure 9, TeleRAG achieves from $1.2\times$ to $1.6\times$ end-to-end speedups for these pipelines, thanks to the combined benefit of reducing CPU workload and utilizing the GPU to perform sorting on similarity distances.
Ablation on cache. Figure 13 illustrates the throughput improvements enabled by caching on the NQ dataset. On average, caching boosts throughput of 2%, 12%, 21%, and 18% with 1, 2, 4, and 8 GPUs, respectively. The benefit on a single GPU is marginal because only a small cache space is used, and the limited cache capacity struggles to accommodate diverse requests. However, as the number of GPUs and overall cache space increases, the cache-aware scheduler effectively assigns micro-batches to the appropriate GPUs to maximize cache overlap, thereby improving prefetching hit rates and yielding significant throughput gains.
Analysis on scheduling overhead. Figure 15 analyzes the benefits and overheads of the prefetching and cache-aware schedulers. It reports the end-to-end latency for a global batch of 128 queries with a mini-batch size of 4 on 4 H200 GPUs, comparing configurations that enable both schedulers, only the prefetching scheduler, or neither. As shown in Figure 15, the prefetching scheduler incurs only a minimal overhead of approximately 37 ms (thus almost invisible in the figure), while the cache-aware scheduler adds a modest overhead of about 220 ms. Overall, both schedulers effectively reduce end-to-end latency in most cases, with minimal additional cost.
6 Related Work
Systems for RAG. Prior works have accelerated RAG by caching the LLM’s KV cache from retrieved documents jin2024ragcache; lu2024turborag; yao2024cacheblend. However, these approaches primarily reduce prefill latency, failing to address the retrieval and decode latencies which often dominate. Other techniques, like speculative retrieval zhang2024accelerating or fine-grained pipelining jiang2024piperag, target different RAG paradigms (e.g., per-token retrieval) and do not apply to the widely-used modular pipelines discussed in §2.1. Unlike these approaches, TeleRAG is the first to tackle the fundamental system challenges of long retrieval latency and large memory requirements for modular RAG with large-scale datastores.
@@ 6 Related Work
Systems for compound LLM applications. Apart from RAG, there is a growing interest in compound or agentic LLM applications, where multiple LLM calls and other applications are combined to serve complex functionalities compound-ai-blog; sigarch; wang2024survey. AI Metropolis xie2024ai accelerates LLM-based multi-agent simulations with out-of-order execution. RAG is a specific type of application in this broader direction, and we propose systems techniques to optimize its execution latency, focusing on the characteristics of retrieval workload.
Vector Index. A separate line of work focuses on optimizing the vector index itself. This includes hardware-specific acceleration (GPU johnson2019billion, FPGA jiang2023co) and hybrid memory-disk systems to scale beyond RAM (DiskANN jayaram2019diskann, SPANN chen2021spann). More recent systems scale graph-based indexes beyond GPU memory khan2024bang; zhang2024fast. These methods often require significant algorithm modifications or remain bottlenecked by CPU–GPU bandwidth. TeleRAG is complementary, optimizing at the system-level by hiding transfer latency within the RAG application’s context, without altering the underlying IVF algorithm.
7 Conclusion
In this paper, we introduced TeleRAG, an inference system that improves RAG pipeline latency and throughput while imposing minimal GPU memory requirements. TeleRAG achieves significant acceleration by employing lookahead retrieval, which hides CPU–GPU data transfer latency by overlapping it with pre-retrieval generation, and by supporting efficient multi-GPU inference through specialized scheduling. Our evaluation shows that TeleRAG significantly improves performance compared to existing state-of-the-art solutions and scales effectively with increased hardware resources.
Appendix A Background: Inverted File Index (IVF) for Vector Search
$\text{k-argmin}_{i=0:N}d(x,y_{i}),$
where $D$ is the dimensionality of the vector determined by the embedding model, and $d$ denotes the distance function, which is typically the L2-norm or inner product johnson2019billion.
To improve search efficiency, the inverted file index (IVF) algorithm is widely used for large-scale vector databases due to its simplicity and effectiveness. IVF partitions the data store into clusters and restricts searches to the most relevant clusters, reducing the computational cost. To obtain the cluster assignment of each stored vector, it performs clustering algorithm such as $k$-means norouzi2013cartesian and partitions the database into $N_{c}$ clusters:
$\{C_{1},C_{2},\ldots,C_{N_{c}}\}=k\text{-means}(Y,N_{c}),$
$\text{Cluster}(y_{i})=\text{argmin}_{j=1:N_{c}}d(y_{i},c_{j}).$
$\{c_{j_{1}},c_{j_{2}},\ldots,c_{j_{L}}\}=L\text{-argmin}_{j=1:C}\ d(x,c_{j}).$
$k\text{-argmin}_{y_{i}\in\cup_{l=1}^{L}C_{j_{l}}}d(x,y_{i}).$
which involves sorting. In this way, IVF reduces search space and accelerates the retrieval process. Here, the hyperparameter $L$ from the first step is also referred as nprobe douze2024faiss. When nprobe is larger, the IVF retrieval will be more accurate as it search more clusters. However, the retrieve latency is longer as more computation and data are needed.
Appendix B Finding Prefetching Amount
$t_{1}=\begin{cases}t_{\text{LLM}},&\text{if }b_{p}\leq B\cdot t_{\text{LLM}},\\[6.0pt] \frac{b_{p}}{B},&\text{if }b_{p}>B\cdot t_{\text{LLM}},\end{cases}$ (1)
As shown in the analysis section, GPU retrieval ($t_{g}$) is generally much faster than CPU retrieval ($t_{c}$). Thus, we assume $t_{c}\gg t_{g}$. As CPU has a limited parallelism, $t_{c}$ usually grows proportionally with the number of clusters to be processed jiang2023co:
$t_{2}=t_{c}=r_{\mathrm{miss}}\times n_{\mathrm{probe}}\times t_{cc},$ (2)
We now examine two cases:
    $b_{p}^{*}=B\cdot t_{\text{LLM}}.$ (3)
    $\frac{\mathrm{d}}{\mathrm{d}b_{p}}(t_{1}+t_{2})=0\implies\frac{1}{B}+\frac{\mathrm{d}r_{\mathrm{miss}}}{\mathrm{d}b_{p}}\cdot n_{\mathrm{probe}}\cdot t_{cc}=0.$ (4)
    From this, we obtain:
    $b_{p}^{*}=B\times n_{\mathrm{probe}}\times t_{cc}\times\Delta r_{\mathrm{miss}},$ (5)
In summary, our analysis shows that the optimal prefetch amount $b_{p}^{*}$ can only lie at one of two points: (1) Prefetch until LLM generation completes (i.e. $b_{p}=B\cdot t_{\text{LLM}}$). (2) A point determined by Eq. 5. However, under typical CPU-GPU bandwidth (e.g., 64 GB/s on PCIe 5), the time spent loading additional clusters often outweighs any retrieval latency reduction from lowering $r_{\mathrm{miss}}$. Consequently, the second scenario in Eq. 5 becomes nearly infeasible in practice. Therefore, on current hardware, prefetching exactly until LLM execution ends is generally the most effective choice.
Appendix C Implementation Details
TeleRAG is implemented in Python and also leverages PyTorch’s pytorch_2 operator ecosystem for efficient computation. The datastore index is initially constructed using FAISS douze2024faiss, and its data structures, such as IVF centroids and cluster data, are converted to PyTorch tensors.
We use separate processes for the schedulers and services in TeleRAG to avoid bottlenecks in Python global interpreter lock. For the communication between the scheduler and the services, we also leverages the ZeroMQ zeromq as the communication framework. The message is encoded and decoded with Pickle in Python. For the LLM sIervice, we use SGLang zheng2024_sglang as the LLM backend. For the retrieval service, to implement the index search with GPU-CPU cooperation, on GPU, we use a single matrix-vector multiplication that computes distances for all prefetched vectors; on CPU, we utilize multithreading in Python to parallelize similarity searches across clusters. We then move the distances computed from CPU to GPU, merge with results on GPU and perform global sorting on GPU.
To monitor the hotness of cached clusters better, we use a decay-based approach to track the hotness of cached clusters. Every cached cluster has a hotness value. Upon fetching to the GPU, the cluster is essentially cached and will be assigned with a initial value $h_{0}=h_{\text{init}}$. Then, when a new round of hotness update starts, the hotness will be divided with a decay factor $d$, and if the cluster is used in the previous round, an hotness update value $h_{\text{inc}}$ will be added the hotness. Therefore, the transition of hotness from the round $r$ to round $r+1$ can be given as:
[…]
In this way, the process can be parallelized easily on GPU, minimize the overhead of hotness tracking in terms of both the time and computation. We also set a cache fraction, which means if the cache exceeds the fraction, we need to evict some clusters. The cache fraction is set to $0.5$ by default. To make evaluation result stable and reproducible, we do cache eviction for the excessive clusters and consolidate the GPU memory after serving each batch. This makes sure that we will not over-evaluate the performance of the caching system.
Both schedulers use greedy search to minimize overhead. The prefetching scheduler groups similar queries into mini-batches, and feeds them into the cache-aware scheduler to assign mini-batches to GPUs. The cache-aware scheduler will get the number of overlapped clusters on each GPU for each mini-batch, and assign the mini-batch to the GPU with the most overlapped clusters.
Appendix D Detailed Index Configurations
Specification Value
Dataset wiki_dpr karpukhin2020dense
Dataset size 2.1 billion tokens
#\# of chunks 21 million
#\# of IVF cluster 4096
Embed model Contriever izacard2021unsupervised
Embed dimension 768
Distance metric Inner Product
Index size 61 GB
Table 4: Detailed configurations of our retrieval index.
Table 4 shows the detailed configurations of our retrieval index used in the evaluation.